OpenStudy (moonlitfate):

Use the Fundamental Theorem of Calculus to determine the answer to the problem; do not use the even/odd properties of integration.

OpenStudy (anonymous):

56665

OpenStudy (moonlitfate):

Problem is: \[\int\limits_{-\frac{ \pi }{ 2 }}^{\frac{ \pi }{ 2 }}(\sin^3x \cos x+\sin x \cos x)dx\]

OpenStudy (anonymous):

\[\sin ^{2} x is meant (\sin(x))^{2}\] so you are using quotient rule for sin square. and cos x anti-derivative was sin x

OpenStudy (anonymous):

Then plug pi/2 and - pi/2 in the equation

OpenStudy (anonymous):

Then take the number when you plug pi/2 subtract the number when you plug -pi/2 and get the answer

OpenStudy (anonymous):

Hope this help

OpenStudy (stamp):

@MoonlitFate find the integral using u substitution (see attachment)

OpenStudy (stamp):

then evaluate the integral from b to a (see attachment 2) verification of answer @ http://www.wolframalpha.com/input/?i=integral+of+sin^3xcosx%2Bsinxcosx+from+-pi%2F2+to+pi%2F2